# A deep backward regression-based scheme for high-dimensional nonlinear partial differential equations

### 16 Mar 2026 v1

@@ 3.2 Error analysis
$\quad+C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\left|R_{i+1,2}^{y}\right|^{2}+\left|R_{i,2}^{z}\right|^{2}+N\left|R_{i+1,1}^{y}\right|^{2}+\left|R_{i,1}^{z}\right|^{2}\right],$
where we use (3.29) and (3.30) in the last inequality. From Young inequality, we know
$\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]\leq 2\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}ds\right]+2h\mathbb{E}\left[\left|\mathcal{V}^{m}_{t_{i}}-\mathcal{Z}_{i}\right|^{2}\right].$ (3.35)
Step 4. By Young inequality, we get
$\mathbb{E}\left[\left|\mathcal{Z}_{i}-\mathcal{V}^{m}_{t_{i}}\right|^{2}\right]\leq$ $2\mathbb{E}\left[\left|\mathcal{Z}_{i}-S_{z,i}\left(p_{i}^{m},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)\right|^{2}\right]+2\mathbb{E}\left[\left|S_{z,i}\left(p^{m}_{i},\Delta W_{i}^{m},X_{i+1}^{\pi,^{m}}\right)-\mathcal{V}^{m}_{t_{i}}\right|^{2}\right].$ (3.36)
In what follows, we handle each term in (3.36) separately.
[…]
$\mathbb{E}\left[\left|\mathcal{Z}_{i}-S_{z,i}\left(p_{i}^{m},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)\right|^{2}\right]=\mathbb{E}\left[\left|R_{i,2}^{z}\right|^{2}\right].$ (3.37)
[…]
$\mathbb{E}\left[\left|S_{z,i}\left(p^{m}_{i},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)-\mathcal{V}^{m}_{t_{i}}\right|^{2}\right]$
$=\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta_{y,i+1}^{*})\frac{(\Delta W_{i}^{m})^{\top}}{h}-\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})\frac{(\Delta W_{i}^{m})^{\top}}{h}\right]\right|^{2}\right]$
$\leq\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta_{y,i+1}^{*})\frac{(\Delta W_{i}^{m})^{\top}}{h}\right|^{2}\right]$
$\leq Nd\mathbb{E}\left[\left|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta_{y,i+1}^{*})\right|^{2}\right]$
$\leq CN+CN\mathbb{E}\left[\left|R_{i+1,2}^{y}\right|^{2}+\left|R_{i+1,1}^{y}\right|^{2}\right],$ (3.38)
where we used (3.28) in the last inequality. Plugging (3.37), (3.38) into (3.36), we derive
$\mathbb{E}\left[\left|\mathcal{Z}_{i}-\mathcal{V}^{m}_{t_{i}}\right|^{2}\right]\leq$ $CN+CN\mathbb{E}\left[\left|R_{i+1,2}^{y}\right|^{2}+h\left|R_{i,2}^{z}\right|^{2}+\left|R_{i+1,1}^{y}\right|^{2}\right].$ (3.39)
Inserting (3.39) into (3.35), we deduce
$\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]\leq 2\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{t_{i}}|^{2}ds\right]+C\mathbb{E}\left[\left|R_{i+1,2}^{y}\right|^{2}+h\left|R_{i,2}^{z}\right|^{2}+\left|R_{i+1,1}^{y}\right|^{2}\right].$ (3.40)
@@
Summing over $i=0,1,\cdots,N-1$, and then inserting (3.34) into the derived inequality, we obtain
$\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]\leq$ $C\mathbb{E}\left[\left|g(X_{T})-g(X_{N}^{\pi,m})\right|^{2}\right]+Ch+C\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]$ (3.41)
$+C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\left|R_{i+1,2}^{y}\right|^{2}+\left|R_{i,2}^{z}\right|^{2}+N\left|R_{i+1,1}^{y}\right|^{2}+\left|R_{i,1}^{z}\right|^{2}\right].$
@@
Adding (3.41) to (3.30), we derive the required error estimate and this ends the proof.
@@ 3.3 Convergence analysis
In this part, we aim to study the convergence rates of the approximate error of the proposed DBR method with the help of the GroupSort deep neural networks (see [1, 17]). Now, let us briefly review it.
Let the set of the GroupSort neural networks is defined as
@@
with $\kappa\in\mathbb{N}^{*},\kappa\geq 2$, be a grouping size, dividing the number of neurons $n_{\ell}=\kappa\tilde{n}_{\ell}$, at each layer $\ell=0,\ldots,L-1$; $\sum\limits_{\ell=0}^{L-1}n_{\ell}$ denotes the width of the network and $L+1$ is the depth; $\zeta_{\kappa}=\left(\zeta_{\kappa}^{\ell}\right)_{\ell=0,\ldots,L-1}$ denotes a specific sequence of activation functions; each nonlinear function $\zeta_{\kappa}^{\ell}$ divides its input into groups of size $\kappa$ and sorts each group in decreasing order; furthermore, by enforcing the parameters of the networks to satisfy, with the Euclidian norm $|\cdot|_{2}$ and the $L_{\infty}$ norm $|\cdot|_{\infty}$,

### 21 May 2026 v3

@@ 3.2 Error analysis
$\quad+C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\varepsilon_{i}^{y}+\varepsilon_{i}^{z}+\frac{N}{K}\right],$
where we use (50) and (51) in the last inequality. From Young inequality, we know
$\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]\leq 2\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{i}|^{2}ds\right]+2h\mathbb{E}\left[\left|\mathcal{V}^{m}_{i}-\mathcal{Z}_{i}\right|^{2}\right].$ (56)
By Young inequality, we get
$\mathbb{E}\left[\left|\mathcal{Z}_{i}-\mathcal{V}^{m}_{i}\right|^{2}\right]\leq$ $2\mathbb{E}\left[\left|\mathcal{Z}_{i}-\widetilde{\mathcal{V}}^{m}_{i}\right|^{2}\right]+2\mathbb{E}\left[\left|\widetilde{\mathcal{V}}^{m}_{i}-\mathcal{V}^{m}_{i}\right|^{2}\right].$ (57)
From (43) and (44), we rewrite (57) as
$\mathbb{E}\left[\left|\mathcal{Z}_{i}-\mathcal{V}^{m}_{i}\right|^{2}\right]\leq$ $C\left(\frac{1}{h}\widetilde{\varepsilon}_{i}^{y}+\widetilde{\varepsilon}_{i}^{z}+\frac{1}{K_{i}h}\right)$ (58)
$\leq$ $C\left(N\widetilde{\varepsilon}_{i}^{y}+\widetilde{\varepsilon}_{i}^{z}+\frac{N}{K_{i}}\right),$
where we used $h=T/N$ and absorbed the constant $T$ into $C$. Inserting (58) into (56), we deduce
$\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]\leq$ $2\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{i}|^{2}ds\right]$ (59)
$+C\left(\widetilde{\varepsilon}_{i}^{y}+h\widetilde{\varepsilon}_{i}^{z}+\frac{1}{K_{i}}\right).$
By (48) and (49), we rewrite (59) as
$\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]\leq$ $2\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{V}^{m}_{i}|^{2}ds\right]$ (60)
$+C\left(\varepsilon_{i}^{y}+h\varepsilon_{i}^{z}+\frac{1}{K_{i}}\right).$
@@
Summing over $i=0,1,\cdots,N-1$, and then inserting (55) into the derived inequality, we obtain
$\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]$ (61)
$+C\sum\limits_{i=0}^{N-1}\left[N\varepsilon_{i}^{y}+\varepsilon_{i}^{z}+\frac{N}{K}\right].$
@@
Adding (61) to (51), we derive the required error estimate and this ends the proof. ∎
Remark 3.1.
Theorem 1 offers a structural advantage over Theorem 4.1 in Hure, Pham, and Warin [29] through a finer decomposition of the error terms on the right-hand side of the inequality. While Theorem 4.1 bounds the approximation error by how well a neural network fits inherently noisy pathwise labels ($\varepsilon_{i}^{\mathcal{N},v}$ and $\varepsilon_{i}^{\mathcal{N},z}$), Theorem 3.1 explicitly isolates the error into two categories: network fitting residuals ($\varepsilon_{i}^{y}$ and $\varepsilon_{i}^{z}$) and statistical Monte Carlo noise ($\frac{1}{K}$). This separation provides a theoretical explanation for the variance-reduction mechanism underlying the DBR method. By taking conditional expectations, the stochastic perturbation $\Delta W_{i}$ is smoothed out prior to loss evaluation. Consequently, the DBR scheme effectively recasts a variance-prone stochastic projection problem into a deterministic regression task, theoretically predicting smoother optimization landscapes, reduced fitting difficulty, and superior generalization capabilities compared to the DBDP method.
@@ 3.3 Convergence analysis
In this section, we investigate the convergence rate of the approximation error for the proposed DBR method utilizing GroupSort deep neural networks (see [1, 17]). To the best of our knowledge, Germain, Pham, and Warin [17] were the first to employ GroupSort deep neural networks to analyze the convergence rates of deep learning methods for PDEs. Building upon their work, we adapt the analytical techniques introduced in Proposition 3.6 of [17] to establish Theorem 3.4. Although the core idea closely parallels theirs, we derive a tailored result specific to our DBR framework. Before proceeding, we briefly review the GroupSort deep neural networks architecture.
The set of the GroupSort neural networks is defined as
@@
with $\kappa\in\mathbb{N}^{*},\kappa\geq 2$, be a grouping size, dividing the number of neurons $n_{\ell}=\kappa\tilde{n}_{\ell}$, at each layer $\ell=0,\ldots,L-1$; $\sum\limits_{\ell=0}^{L-1}n_{\ell}$ denotes the width of the network and $L+1$ is the depth; $\zeta_{\kappa}=\left(\zeta_{\kappa}^{\ell}\right)_{\ell=0,\ldots,L-1}$ denotes a specific sequence of activation functions; each nonlinear function $\zeta_{\kappa}^{\ell}$ divides its input into groups of size $\kappa$ and sorts each group in decreasing order; furthermore, by enforcing the parameters of the networks to satisfy, with the Euclidean norm $|\cdot|_{2}$ and the $L_{\infty}$ norm $|\cdot|_{\infty}$,